OpenStudy (anonymous):

how do u find 3 points for this equation 3x+2y=2?

OpenStudy (anonymous):

substitute 3 values for x .......solve for Y....

OpenStudy (anonymous):

guesss whooooo??? xDD

OpenStudy (anonymous):

yay thankss :P

OpenStudy (phi):

do you know how to "solve for y"? that means put y on the left side by itself.

OpenStudy (anonymous):

i keep getting fractions and i cant graph fractions

OpenStudy (anonymous):

3x+2y=2 2y=-3x+2 y=(-3x+2)/2 just plug in X and you will get Y I recommend you use a table.........thats what that one dude said just ask for an answer.

OpenStudy (anonymous):

kayy

OpenStudy (phi):

if you start with y= (-3/2)x +1 try x=0 x=2 x=-2

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

notice that x=2 will "cancel" the 2 in the denominator of -3/2, so no fraction